OpenStudy (anonymous):

Determine the width of the patio. Omar must put 50 feet of fencing around his rectangular patio. The length of the patio is one foot longer than three times the width. What is the solution set for 2w + 2(3w + 1) = 50 given the replacement set {6, 7, 8, 9}?

OpenStudy (jjuden):

width =6 feet, length = 19 feet

OpenStudy (jjuden):

2w + 2(3w + 1) = 50 2w + 6w + 2 = 50 8w + 2 = 50 8w = 48 w = 6 then 6*3+1=L 18+1=L 19=L